# FedRecon: Missing Modality Reconstruction in Heterogeneous Distributed Environments

### v2

@@ Abstract.
Multimodal data are often incomplete and exhibit Non-Independent and Identically Distributed (Non-IID) characteristics in real-world scenarios. These inherent limitations lead to both modality heterogeneity through partial modality absence and data heterogeneity from distribution divergence, creating fundamental challenges for effective federated learning (FL). To address these coupled challenges, we propose FedRecon, the first method targeting simultaneous missing modality reconstruction and Non-IID adaptation in multimodal FL. Our approach first employs a lightweight Multimodal Variational Autoencoder (MVAE) to reconstruct missing modalities while preserving cross-modal consistency. Distinct from conventional imputation methods, we achieve sample-level alignment through a novel distribution mapping mechanism that guarantees both data consistency and completeness. Additionally, we introduce a strategy employing global generator freezing to prevent catastrophic forgetting, which in turn mitigates Non-IID fluctuations. Extensive evaluations on multimodal datasets demonstrate FedRecon’s superior performance in modality reconstruction under Non-IID conditions, surpassing state-of-the-art methods. Our code has been released on https://github.com/Wings-Of-Disaster/FedRecon.
Multimodal Federated Learning, Multimodal Variational Autoencoder, Modality Reconstruction
*Corresponding author.
ccs: Computing methodologies Distributed artificial intelligence
@@
Figure 1. Impact of modality and data heterogeneity on the convergence of the global model in MFL. Client 1 and 2 have complete modalities but suffer from data heterogeneity due to divergent local distributions. In contrast, client $i$ and client $N$ are missing modalities and perform unimodal self-reconstruction and cross-modal reconstruction respectively.
[…]
This figure illustrates the impact of data and modality heterogeneity on the convergence of the global model in MFL. First, data heterogeneity arises from non-identical data distributions across clients, which leads to model updates that are biased or divergent, hindering stable convergence. Second, due to the presence of missing modalities, reconstruction is often required. However, such reconstruction can suffer from misalignment at the label level or at the sample level. These alignment issues degrade the quality of the reconstructed modalities, further impairing the convergence of the global model.
1. Introduction
@@
Federated Learning (FL) has gained significant attention as a promising framework for decentralized machine learning, ensuring privacy protection in distributed environments (Li et al., 2020a, c; Qin et al., 2023; Xiong et al., 2024; Liu et al., 2025a). While existing FL methods demonstrate strong performance in unimodal settings (Chen et al., 2025; Yu et al., 2024b), their inability to capture cross-modal interactions fundamentally limits real-world applicability (Huang et al., 2021, 2024a; Lin et al., 2023). To bridge this gap, Multimodal Federated Learning (MFL) has emerged as a critical solution by enabling complementary information fusion across modalities (Zhao et al., 2022; Che et al., 2023), with successful deployments in autonomous driving (Cui et al., 2019), video action capture (Ofli et al., 2013), and multimedia sentiment analysis (Kaur and Kautish, 2019; Liu et al., 2025b). By explicitly modeling inter-modal dependencies, MFL achieves holistic understanding of complex phenomena (Li et al., 2024; Chen and Zhang, 2024), enabling precise real-world inference under data scarcity.
@@
The majority of existing MFL approaches presume full modality availability across clients (Huang et al., 2024b). However, this assumption is often misaligned with real-world conditions, where data availability is frequently restricted. For instance, certain clients may possess limited resources or only support unimodal data acquisition, whereas others may have access to more advanced multimodal setups (Poria et al., 2017). Furthermore, the quality of the data can be severely impacted by fluctuations during the collection process, leading to inconsistencies that undermine the reliability of the captured information (Hassan et al., 2021). In this study, we formally consider a MFL system with $K$ clients and up to $|M|$ modalities. All clients and the central server share a common model architecture, but each client $i$ has access to a subset $M_{i}\subseteq M$ of the available modalities. This results in a scenario where $M_{i}\neq M_{i^{\prime}}$ for some clients $i,i^{\prime}\in[K]$. Consequently, the system introduces data heterogeneity across clients and disrupts inter-modal connections within each client, exacerbating the challenges in MFL.
@@
The global model $F$ requires comprehensive processing of all modalities, precluding the discarding of any modality during local training. This necessitates reconstruction of missing modalities $M^{\prime}_{i}=M\setminus M_{i}$ through feature generation from available client data. Existing reconstruction methods range from naive zero-filling (Van Buuren and Van Buuren, 2012) to advanced self-supervised synthesis (Li et al., 2017; Zong et al., 2024). However, zero-filling frequently degrades model accuracy, while self-supervised methods that reduce label dependency fail to adequately address the complex challenge of cross-modal alignment (Ma et al., 2021). Considering the limited storage and communication capacity of clients in MFL, it is crucial to adopt a lightweight and more efficient alignment method.
@@
While some methods have focused on the alignment issues during the modality reconstruction process (Xiong et al., 2023; Yu et al., 2023; Wang et al., 2024; Yan et al., 2024), they inadequately preserve semantic consistency across samples sharing identical labels. As demonstrated in Figure 1, reconstructed modalities show inconsistent feature representations despite matching labels. This arises because existing methods align modalities through label correspondence but ignore semantic nuances specific to individual samples. Such limitations highlight that label agreement alone cannot guarantee multimodal coherence, as inherent variations within classes persist across clients.
@@
In this paper, we present FedRecon, a novel framework designed to address missing modality reconstruction in distributed heterogeneous environments. By employing a lightweight Multimodal Variational Autoencoder (MVAE) (Kingma et al., 2014; Wu and Goodman, 2018; Shi et al., 2019; Palumbo et al., 2023) as the core engine, FedRecon performs cross-modal alignment and modality imputation while overcoming client resource constraints in training and communication. Compared to computationally intensive alternatives like GANs (Goodfellow et al., 2014) and Diffusion models (Ho et al., 2020), our architecture substantially reduces training time and communication cost through its streamlined encoder-decoder design, with complete efficiency investigation provided in Appendix A. Critically, we introduce a latent vector mapping mechanism that uniquely establishes sample-level alignment through direct feature coupling across modalities, ensuring semantic consistency during reconstruction. Furthermore, we maintain a frozen auxiliary model initialized from the global MVAE to stabilize the generation process in Non-IID environments, collaborating with local models to generate missing modalities.
@@
  • To the best of out knowledge, we are the first to leverage MVAEs for partial modality reconstruction in MFL, addressing scenarios with variable missing ratios while reducing training and communication costs.
@@
  • Experimental results validate state-of-the-art performance, surpassing baseline methods by 2.99% and 4.57% absolute gains on MELD and CrisisMMD benchmarks respectively, confirming the efficacy of FedRecon in real-world scenarios.
2. Related Work
2.1. Multimodal Variational Autoencoder
The Multimodal Variational Autoencoder (MVAE) extends the framework of Variational Autoencoders (VAE) (Kingma et al., 2014) to enable joint learning and generation across heterogeneous data modalities. Wu and Goodman (Wu and Goodman, 2018) pioneered this direction by introducing a scalable architecture capable of learning shared representations in weakly supervised settings, where only a subset of data samples contain all modalities. Subsequent work further advanced the generative coherence and fidelity of MVAEs. Shi et al. (Shi et al., 2019) proposed a mixture-of-experts formulation (MMVAE) that decomposes latent spaces into shared and modality-specific subspaces while ensuring coherent joint generation through multimodal integration. Palumbo et al. (Palumbo et al., 2023) enhanced this framework by explicitly isolating distributions of shared and private latent factors, thereby improving robustness to variations in latent space design without compromising semantic consistency. Further innovations integrated Diffusion models into the MVAE architecture (Palumbo et al., 2024), demonstrating enhanced unconditional generation performance for complex real-world data modalities. Despite their effectiveness in aligning modalities through joint latent space learning, current methods exhibit limitations in maintaining consistent cross-modal alignment across diverse samples sharing the same label, especially when used on highly variable datasets or data with real-world noise. Notably, FedRecon mitigates these deficiencies through a simple mapping mechanism, demonstrating practical efficacy in cross-modal alignment.
2.2. Missing Modality Reconstruction in MFL
Reconstructing missing modalities in MFL ensures robust model performance under heterogeneous data distributions. Xiong et al. propose CACMRN (Xiong et al., 2023), a client-adaptive transformer leveraging instance relationships and federated optimization to prevent local overfitting during cross-modal reconstruction. Wang et al. develop FedMMR (Wang et al., 2024), aligning feature spaces via modality synthesis while integrating reconstructed and existing modalities as complementary inputs. Liu et al. design FedMobile (Liu et al., 2025c), prioritizing cross-node latent feature sharing to address extreme modality incompleteness. Yu et al. introduce FedInMM (Yu et al., 2024a), utilizing LSTM-based dynamic weighting and temporal fusion for irregular multimodal data. While effective in reconstructing modalities under limited samples in MFL, these methods overlook sample-level alignment across clients with identical labels. Additionally, they inadequately address instability caused by Non-IID client data during cross-modal generation. FedRecon stabilizes training by freezing a global generator replica to reduce distributional shifts, thereby mitigating fluctuations induced by client-specific data divergence.
3. Problem Formulation
@@
(1) $\mathcal{L}_{i}(X_{i},y_{i};\phi_{i})=\mathcal{L}_{\text{task}}\left(H\left(\oplus_{m\in M_{i}}\{f_{i,m}\}\right),y_{i}\right),$
where $\phi_{i}$ denotes the set of parameters for client $i$, and $f_{i,m}=E_{i,m}(X_{i,m})$ represents the features extracted by the encoder $E_{i,m}$ for modality $m$. Here, $\oplus$ refers to the fusion operation that combines the features from all available modalities into a unified representation. In line with multi-task MFL frameworks (Wang et al., 2024; Feng et al., 2023), the global objective $\mathcal{L}$ for MFL is expressed as:
(2) $\min_{\phi}\mathcal{L}(F(\phi))=\frac{1}{K}\sum_{i=1}^{K}\mathcal{L}_{i}(F_{i}(X_{i},y_{i};\phi_{i})),$
where $\mathcal{L}_{i}$ denotes the local loss function for client $i$, $F_{i}$ represents the client-specific model trained on its available modalities, and $\phi$ refers to the parameters of the global model.
4. Proposed Method
Figure 2. (a) The workflow of FedRecon, illustrating the process of reconstructing missing modality in distributed heterogeneous environments. (b) The communication mechanism of FedRecon, where each client downloads and submits model according to its modalities, reducing overhead during distributed learning.
A promising approach to mitigate the missing modality issue is to enforce consistency across clients (Wang et al., 2024). Our method builds on this foundation with two key components: (1) a lightweight MVAE that reconstructs missing modalities via an innovative distribution mapping mechanism to preserve cross-modal consistency, and (2) a global generator freezing strategy to stabilize Non-IID learning dynamics. The complete workflow is shown in Figure 2.
4.1. Missing Modality Reconstruction
Considering the challenges of training cost, storage capacity, and communication overhead in MFL scenarios, we conducted a comprehensive survey of various generative architectures. Among them, VAE emerges as particularly advantageous due to its relatively compact parameter structure. This characteristic translates to diminished computational requirements and storage consumption on client devices, while simultaneously minimizing data transmission burdens. Such attributes render VAE exceptionally suitable for MFL environments where efficient resource utilization is critical.
@@ 4.1. Missing Modality Reconstruction
The goal of VAE is to maximize the log-likelihood of the data, but due to computational complexity, it is often optimized via a lower bound (Kingma et al., 2014). The optimization objective of VAE is:
(3) $\mathcal{L}_{\text{G}_{i,m}}(X_{i,m})=-\mathbb{E}_{q_{\psi}(z|X_{i,m})}\left[\log\frac{p_{\theta}(X_{i,m},z)}{q_{\psi}(z|X_{i,m})}\right]$
@@
where the input data $X_{i,m}$ corresponds to modality $m$, with $z$ denoting the latent variables. The approximate posterior $q_{\psi}(z|X_{i,m})$ is parameterized by the encoder with learnable parameters $\psi$, and $p_{\theta}(X_{i,m},z)=p(z)p_{\theta}(X_{i,m}|z)$ is the joint distribution generated by the decoder with parameters $\theta$.
As a result, the loss can be divided into two components. The reconstruction loss $\mathcal{L}_{\text{recon}}$ quantifies the ability of the decoder to accurately reconstruct the input data $X_{i,m}$ based on the latent variable $z$, whereas the KL divergence term $\mathcal{L}_{\text{KL}}$ imposes a regularization on the posterior distribution $q_{\psi}(z|X_{i,m})$, encouraging it to approximate the prior distribution $p(z)$. Notably, the VAE-based generator we adopt is label-agnostic. By leveraging the label space $Y_{i}$, the generator models a distribution that aligns with the client’s attributes, serving as a high-level abstraction of the raw data and mitigating privacy risks without exposing sensitive information.
As mentioned above, we perform self-reconstruction for each modality to obtain highly correlated data on missing modalities. Our generator is particularly suited for MFL environments, producing high-quality data while maintaining a lightweight architecture.
4.2. Cross-modal Alignment and Modality Awareness Fusion
@@
While advanced MVAE architectures leverage shared latent variables (Shi et al., 2019; Palumbo et al., 2023) to enable cross-modal interaction, they often suffer from limited diversity in reconstructed outputs and fail to achieve precise sample-level alignment across modalities. Our framework introduces a distribution re-encoding mechanism through mapping modelsl $T_{n,m}$, which adapt the latent variable $z_{m}$ to better suit each target modality. This design provides modality-specific latent inputs for each decoder $G_{i,m}^{\text{dec}}$, allowing for both decoupled distribution modeling and more accurate sample-level alignment across modalities. The optimization objective in the multimodal setting becomes:
(4) $\mathcal{L}_{\text{G}_{i,m}}(X_{i})=-\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}_{q_{\psi_{m}}(z_{m}|X_{i,m})}\left[\log\frac{p_{\Theta}(X_{i},z_{m})}{q_{\psi_{m}}(z_{m}|X_{i,m})}\right],$
@@
Compared to optimizing feature alignment across different modalities using similarity at the feature level (Wang et al., 2024; Liu et al., 2023), directly aligning a compressed distribution proves to be more efficient. The structure of the model $T_{n,m}$ is based on a simple MLP architecture, which projects the latent distribution from modality $m$ onto that of modality $n$ in the latent space. However, we found that directly training the model using the objective in Equation 4 leads to instability and poor convergence. To address this, we decompose the training process into two stages. In the first stage, we jointly train the encoder and decoder components using the objective in Equation 3 to learn modality-specific representations and enable accurate reconstruction. In the second stage, we freeze the encoder and decoder parameters and exclusively optimize the mapping models $T_{n,m}$ so that the latent variable encoded from modality $m$ can be transformed to faithfully reconstruct the data from modality $n$.
To further stabilize this training process, we optionally apply KL divergence as a regularization term to promote convergence when the mapping $T_{n,m}$ is hard to optimize. This encourages the mapped distribution to align better with the target latent space. The final training objective for $T_{n,m}$ is defined as:
(5) $\mathcal{L}_{\text{match}}=\alpha\cdot D_{\text{KL}}\left(T_{n,m}(z_{m})\,\|\,z_{n}\right)+\mathcal{L}_{\text{recon}},$
where $\alpha\in\{0,1\}$ is a binary coefficient controlling the use of the KL regularization term, and $\mathcal{L}_{\text{recon}}$ denotes the reconstruction loss.
To enhance semantic consistency between the reconstructed and available modalities, we adopt an attention-based feature fusion mechanism inspired by (Feng et al., 2023). Given the extracted features $f_{i,m}$ and its reconstructed counterpart $f_{i,m}^{\prime}$, we concatenate them into a joint representation $h=[f_{i,m};f_{i,m}^{\prime}]$, and compute attention weights as follows:
(6) $u=\tanh(Wh+$ $b),\quad a=\text{softmax}(u^{T}c),$
$f_{i,m}^{*}$ $=\sum_{i}a_{i}h_{i},$
@@
4.3. Global Generator Freezing Strategy
In Section 4.2, we train the generator using complete modality data $X^{j}_{i}$, and employ it during inference to predict the missing modality $X^{k}_{i,m^{\prime}}$ from the available modality $X^{k}_{i,m}$. However, in Non-IID settings, the missing modality $X^{k}_{i,m^{\prime}}$ may exhibit a noticeable distribution shift with respect to the local data of client $i$, making it underrepresented during local training and thus prone to forgetting. In such cases, the missing data $X^{k}_{i,m^{\prime}}$ may be better aligned with the distribution of another client $i^{\prime}$, where similar samples $X^{j}_{i^{\prime}}$ are more common. Consequently, relying solely on local generators can lead to suboptimal reconstruction performance.
Figure 3. An illustrative example of the Global Generator Freezing Strategy on the PolyMNIST dataset.
To mitigate this, we propose a global generator freezing strategy, where each client maintains two generators: one global generator initialized at deployment and kept frozen, and another updated locally. This allows clients to fall back on the global generator when their own reconstruction suffers from data shift or forgetting.
As illustrated in Figure 3, we replace the original synthetic data with a concatenation of outputs from both global and local generators, and reduce the weight of synthetic features during the attention fusion process to achieve a balancing effect. Despite the additional requirement of storing both generators and mapping models on each client, our approach remains significantly more lightweight than GAN- or Diffusion-based alternatives. The detailed algorithmic workflow is outlined in Algorithm 1.
5. Experiments
5.1. Experimental Setups
Datasets. To evaluate the effectiveness of our MFL framework, we utilize two real-world multimodal datasets: (1) MELD (Poria et al., 2018), a multiparty conversational dataset comprising over 9k utterances with both audio and textual transcripts from the TV series Friends, and (2) CrisisMMD (Alam et al., 2018), which contains 18.1k tweets accompanied by both visual and textual information, collected during various crisis events. To further assess the generative quality of our MVAE model, we additionally adopt two benchmark multimodal datasets: (3) PolyMNIST (Sutter et al., 2021a), a synthetic dataset featuring five modalities, where each consists of MNIST digits composited over diverse background images; and (4) CUB (Wah et al., 2011), the Caltech-UCSD Birds dataset, which pairs bird images with fine-grained textual descriptions, posing a challenging setting due to the modality-specific nuances in visual and textual representations.
Algorithm 1 Two-Stage Federated Training for FedRecon
1:for round $t=1,2,\dots,T$ do
2:     for client $k=1$ to $K$ in parallel do
3:         Stage 1: Local training of MVAE
5:         $\widetilde{T}_{k}^{(t)},T_{k}^{(t)}\leftarrow T^{(t-1)}$
7:         Stage 2: Freeze MVAE and train mapping modules
8:         $\{G_{k,m}^{(t)}\}_{m=1}^{M}\leftarrow\text{Freeze}(\{G_{k,m}^{(t)}\}_{m=1}^{M})$
9:         Update $T_{k}^{(t)}\leftarrow\text{SGD}(T_{k}^{(t)})$ for $L_{2}$ epochs
10:         Stage 3: Update task model
13:     end for
14:     $\{G_{m}^{(t)}\}_{m=1}^{M}\leftarrow\left\{\frac{1}{K}\sum_{k=1}^{K}G_{k,m}^{(t)}\right\}_{m=1}^{M}$
15:     $T^{(t)}\leftarrow\frac{1}{K}\sum_{k=1}^{K}T_{k}^{(t)}$
16:     $\phi^{(t)}\leftarrow\frac{1}{K}\sum_{k=1}^{K}\phi_{k}^{(t)}$
17:end for

### v3

@@ Abstract
Multimodal data are often incomplete and exhibit Non-Independent and Identically Distributed (Non-IID) characteristics in real-world scenarios. These inherent limitations lead to both modality heterogeneity through partial modality absence and data heterogeneity from distribution divergence, creating fundamental challenges for effective federated learning (FL). To address these coupled challenges, we propose FedRecon, the first method targeting simultaneous missing modality reconstruction and Non-IID adaptation in multimodal FL. Our approach first employs a lightweight Multimodal Variational Autoencoder (MVAE) to reconstruct missing modalities while preserving cross-modal consistency. Distinct from conventional imputation methods, we achieve sample-level alignment through a novel distribution mapping mechanism that guarantees both data consistency and completeness. Additionally, we introduce a strategy employing global generator freezing to prevent catastrophic forgetting, which in turn mitigates Non-IID fluctuations. Extensive evaluations on multimodal datasets demonstrate FedRecon’s superior performance in modality reconstruction under Non-IID conditions, surpassing state-of-the-art methods. The code will be released upon paper acceptance.
@@
Figure 1: Impact of modality and data heterogeneity on the convergence of the global model in MFL. Client 1 and 2 have complete modalities but suffer from data heterogeneity due to divergent local distributions. In contrast, client $i$ and client NN are missing modalities and perform unimodal self-reconstruction and cross-modal reconstruction respectively.
Introduction
@@
Federated Learning (FL) has gained significant attention as a promising framework for decentralized machine learning, ensuring privacy protection in distributed environments (Li et al. 2020a, b; Qin, Li, and Zhu 2023; Xiong et al. 2024; Liu et al. 2025a). While existing FL methods demonstrate strong performance in unimodal settings (Chen et al. 2025; Yu et al. 2024a), their inability to capture cross-modal interactions fundamentally limits real-world applicability (Huang et al. 2021, 2024a; Lin et al. 2023). To bridge this gap, Multimodal Federated Learning (MFL) has emerged as a critical solution by enabling complementary information fusion across modalities (Zhao, Barnaghi, and Haddadi 2022; Che et al. 2023), with successful deployments in autonomous driving (Cui et al. 2019), video action capture (Ofli et al. 2013), and multimedia sentiment analysis (Kaur and Kautish 2019). By explicitly modeling inter-modal dependencies, MFL achieves holistic understanding of complex phenomena (Li et al. 2024; Chen and Zhang 2024), enabling precise real-world inference under data scarcity.
@@
The majority of existing MFL approaches presume full modality availability across clients (Huang et al. 2024b; Gao et al. 2025). However, this assumption is often misaligned with real-world conditions, where data availability is frequently restricted. For instance, certain clients may possess limited resources or only support unimodal data acquisition, whereas others may have access to more advanced multimodal setups (Poria et al. 2017; Liu et al. 2025b). Furthermore, the quality of the data can be severely impacted by fluctuations during the collection process, leading to inconsistencies that undermine the reliability of the captured information (Hassan, Leong, and Schneider 2021). In this study, we formally consider a MFL system with $K$ clients and up to $|M|$ modalities. All clients and the central server share a common model architecture, but each client $i$ has access to a subset $M_{i}\subseteq M$ of the available modalities. This results in a scenario where $M_{i}\neq M_{i^{\prime}}$ for some clients $i,i^{\prime}\in[K]$. Consequently, the system introduces data heterogeneity across clients and disrupts inter-modal connections within each client, exacerbating the challenges in MFL.
@@
The global model $F$ requires comprehensive processing of all modalities, precluding the discarding of any modality during local training. This necessitates reconstruction of missing modalities $M^{\prime}_{i}=M\setminus M_{i}$ through feature generation from available client data. Existing reconstruction methods range from naive zero-filling (Van Buuren and Van Buuren 2012) to advanced self-supervised synthesis (Li, Hu, and Nie 2017; Zong, Aodha, and Hospedales 2024). However, zero-filling frequently degrades model accuracy, while self-supervised methods that reduce label dependency fail to adequately address the complex challenge of cross-modal alignment (Ma et al. 2021). Considering the limited storage and communication capacity of clients in MFL, it is crucial to adopt a lightweight and more efficient alignment method.
@@
While some methods have focused on the alignment issues during the modality reconstruction process (Xiong et al. 2023; Yu et al. 2023; Wang et al. 2024; Yan et al. 2024), they inadequately preserve semantic consistency across samples sharing identical labels. As demonstrated in Figure 1, reconstructed modalities show inconsistent feature representations despite matching labels. This arises because existing methods align modalities through label correspondence but ignore semantic nuances specific to individual samples. Such limitations highlight that label agreement alone cannot guarantee multimodal coherence, as inherent variations within classes persist across clients [^1] Related work is discussed in Appendix A..
@@
In this paper, we present FedRecon, a novel framework designed to address missing modality reconstruction in heterogeneous distributed environments. By employing a lightweight Multimodal Variational Autoencoder (MVAE) (Kingma, Welling et al. 2014; Wu and Goodman 2018; Shi et al. 2019; Palumbo, Daunhawer, and Vogt 2023) as the core engine, FedRecon performs cross-modal alignment and modality imputation while overcoming client resource constraints in training and communication. Compared to computationally intensive alternatives like GANs (Goodfellow et al. 2014) and Diffusion models (Ho, Jain, and Abbeel 2020), our architecture substantially reduces training time and communication cost through its streamlined encoder-decoder design, with complete efficiency investigation provided in Appendix B. Critically, we introduce a latent vector mapping mechanism that uniquely establishes sample-level alignment through direct feature coupling across modalities, ensuring semantic consistency during reconstruction. Furthermore, we maintain a frozen auxiliary model initialized from the global MVAE to stabilize the generation process in Non-IID environments, collaborating with local models to generate missing modalities.
@@
  • We are the first to apply MVAEs for realistic partial modality reconstruction in MFL under the setting where modality absence is randomly simulated based on missing rates, reflecting real-world scenarios.
@@
  • Experimental results validate state-of-the-art performance, surpassing baseline methods by 3.06% and 5.28% absolute gains on MELD and CrisisMMD benchmarks. Quality evaluations on CUB and PolyMNIST further demonstrate the effectiveness of FedRecon.
Problem Formulation
@@
$\mathcal{L}_{i}(X_{i},y_{i};\phi_{i})=\mathcal{L}_{\text{task}}\left(H\left(\oplus_{m\in M_{i}}\{f_{i,m}\}\right),y_{i}\right),$ (1)
where $f_{i,m}=E_{i,m}(X_{i,m})$ is the feature extracted by encoder $E_{i,m}$ for modality $m$, and $\oplus$ denotes the fusion of available modalities. Following multi-task MFL frameworks (Wang et al. 2024; Feng et al. 2023), the global objective is:
$\min_{\phi}\mathcal{L}(F(\phi))=\frac{1}{K}\sum_{i=1}^{K}\mathcal{L}_{i}(F_{i}(X_{i},y_{i};\phi_{i})),$ (2)
where $\phi_{i}$ denotes the set of parameters for client $i$ and $F_{i}$ represents the client-specific model trained on $M_{i}$.
Proposed Method
Figure 2: (a) The workflow of FedRecon, illustrating the process of reconstructing missing modality in distributed heterogeneous environments. (b) The communication mechanism of FedRecon, where each client downloads and submits model according to its modalities, reducing overhead during distributed learning.
A promising approach to mitigate the missing modality issue is to enforce consistency across clients (Wang et al. 2024). Our method builds on this foundation with two key components: (1) a lightweight MVAE that reconstructs missing modalities via an innovative distribution mapping mechanism to preserve cross-modal consistency, and (2) a global generator freezing strategy to stabilize Non-IID learning dynamics. The complete workflow is shown in Figure 2.
Missing Modality Reconstruction
Considering the challenges of training cost, storage capacity, and communication overhead in MFL scenarios, we conducted a comprehensive survey of various generative architectures. Among them, VAE emerges as particularly advantageous due to its relatively compact parameter structure. This characteristic translates to diminished computational requirements and storage consumption on client devices, while simultaneously minimizing data transmission burdens. Such attributes render VAE exceptionally suitable for MFL where efficient resource utilization is critical.
@@
The goal of VAE is to maximize the log-likelihood of the data, but due to computational complexity, it is often optimized via a lower bound (Kingma, Welling et al. 2014). The optimization objective of VAE is:
$\mathcal{L}_{\text{G}_{i,m}}(X_{i,m})=-\mathbb{E}_{q_{\psi}(z|X_{i,m})}\left[\log\frac{p_{\theta}(X_{i,m},z)}{q_{\psi}(z|X_{i,m})}\right]$ (3)
@@
where the input data $X_{i,m}$ corresponds to modality $m$, with $z$ denoting the latent variables. The approximate posterior $q_{\psi}(z|X_{i,m})$ is parameterized by $G_{i,m}^{\text{enc}}$ with learnable parameters $\psi$, and $p_{\theta}(X_{i,m},z)=p(z)p_{\theta}(X_{i,m}|z)$ is the joint distribution generated by $G_{i,m}^{\text{dec}}$ with parameters $\theta$.
As a result, the loss can be divided into two components. The reconstruction loss $\mathcal{L}_{\text{recon}}$ quantifies the ability of the decoder to accurately reconstruct the input data $X_{i,m}$ based on the latent variable $z$, whereas the KL divergence term $\mathcal{L}_{\text{KL}}$ regularizes the posterior distribution $q_{\psi}(z|X_{i,m})$, encouraging it to approximate the prior distribution $p(z)$. Notably, the VAE-based generator we adopt is label-agnostic. By leveraging the label space $Y_{i}$, the generator models a distribution that aligns with the client’s attributes, serving as a high-level abstraction of the raw data and mitigating privacy risks without exposing sensitive information.
Cross-modal Alignment and Modality Awareness Fusion
@@
While advanced MVAE architectures leverage shared latent variables (Shi et al. 2019; Palumbo, Daunhawer, and Vogt 2023) to enable cross-modal interaction, they often suffer from limited diversity in reconstructed outputs and fail to achieve precise sample-level alignment across modalities. Our framework introduces a distribution re-encoding mechanism through mapping models $T_{n,m}$, which adapt the latent variable $z_{m}$ to better suit each target modality $n$. This design provides modality-specific latent inputs for each decoder, allowing for both decoupled distribution modeling and more accurate sample-level alignment across modalities. The optimization objective in the multimodal setting becomes:
$\mathcal{L}_{\text{G}_{i,m}}(X_{i})=-\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}_{q_{\psi_{m}}(z_{m}|X_{i,m})}\left[\log\frac{p_{\Theta}(X_{i},z_{m})}{q_{\psi_{m}}(z_{m}|X_{i,m})}\right],$ (4)
@@
Compared to optimizing feature alignment across different modalities using similarity at the feature level (Wang et al. 2024; Liu et al. 2023), directly aligning a compressed distribution proves to be more efficient. The structure of the model $T_{n,m}$ is based on a simple MLP architecture, which projects the latent distribution from modality $m$ onto that of modality $n$ in the latent space. However, we found that directly training the model using the objective in Equation 4 leads to instability and poor convergence. To address this, we decompose the training process into two stages. In the first stage, we jointly train the encoder and decoder components using the objective in Equation 3 to learn modality-specific representations and enable accurate reconstruction. In the second stage, we freeze the encoder and decoder parameters and exclusively optimize $T_{n,m}$ so that the latent variable encoded from modality $m$ can be transformed to faithfully reconstruct the data from modality $n$.
To further stabilize this training process, we optionally apply KL divergence as a regularization term to promote convergence when $T_{n,m}$ is hard to optimize. This encourages the mapped distribution to align better with the target latent space. The final training objective for $T_{n,m}$ is defined as:
$\mathcal{L}_{\text{match}}=\gamma\cdot D_{\text{KL}}\left(T_{n,m}(z_{m})\,\|\,z_{n}\right)+\mathcal{L}_{\text{recon}},$ (5)
where $\gamma\in\{0,1\}$ is a binary coefficient controlling the use of the KL regularization term.
To enhance semantic consistency between the reconstructed and available modalities, we adopt an attention-based feature fusion mechanism inspired by (Feng et al. 2023). Given the extracted features $f_{i,m}$ and its reconstructed counterpart $f_{i,m}^{\prime}$, we concatenate them into a joint representation $h=[f_{i,m};f_{i,m}^{\prime}]$, and compute attention weights as follows:
$u=\tanh(Wh+$ $b),\quad a=\text{softmax}(u^{T}c),$ (6)
[…]
@@
Global Generator Freezing Strategy
In Section Cross-modal Alignment and Modality Awareness Fusion, we train the generator using complete modality data $X^{j}_{i}$, and employ it during inference to predict the missing modality $X^{k}_{i,m^{\prime}}$ from the available modality $X^{k}_{i,m}$. However, in Non-IID settings, the missing modality $X^{k}_{i,m^{\prime}}$ may exhibit a noticeable distribution shift with respect to the local data of client $i$, making it underrepresented during local training and thus prone to forgetting. In such cases, the missing data $X^{k}_{i,m^{\prime}}$ may be better aligned with the distribution of another client $i^{\prime}$, where similar samples $X^{j}_{i^{\prime}}$ are more common. Consequently, relying solely on local generators can lead to suboptimal reconstruction performance.
Figure 3: An illustrative example of the Global Generator Freezing Strategy on the PolyMNIST dataset.
To mitigate this, we propose a global generator freezing strategy, where each client maintains two generators: one global generator initialized at deployment and kept frozen, and another updated locally. This allows clients to fall back on the global generator when their own reconstruction suffers from data shift or forgetting. As illustrated in Figure 3, we replace the original synthetic data with a concatenation of outputs from both global and local generators, and reduce the weight of synthetic features during the attention fusion process to achieve a balancing effect. Despite the additional requirement of storing both generators and mapping models on each client, our approach remains significantly more lightweight than GAN- or Diffusion-based alternatives. The detailed algorithmic workflow is outlined in Appendix C.
Experiments
Experimental Setups